### v1

@@
Coupled Continuous-Discontinuous Galerkin Finite Element Solver for Compound Flood Simulations
Chayanon Wichitrnithed namo@utexas.edu Eirik Valseth Shintaro Bunya Ethan J. Kubatko Clint Dawson
@@ 1 Introduction
Numerical models based on finite element (FE) discretizations to simulate hydrodynamics in coasts, estuaries, and the ocean have widespread use in engineering, marine, and other scientific disciplines. The model that is the focus of our current investigation are the depth-averaged two dimensional Shallow Water Equations (SWE) [1]. In this particular case, we use the SWE to model flows as a result of the interaction between rivers, rainfall runoff, and storm surge, i.e., compound floods [2]. Famous examples from the last decade includes Hurricane Harvey (2017) [3] Hurricane Irma (2017) [4], and Hurricane Florence (2018) [5]. The flow resulting from the interaction of these processes requires careful treatment to ensure accurate modeling and has therefore been frequently studied in recent literature, see, e.g., [6, 7, 8, 9]. The aim of the present paper is to introduce a new numerical model in which we use a modified form of the SWE to develop to simulate compound flooding during tropical cyclones [2].
@@
A widely used existing FE based numerical model for the SWE is the ADvanced CIRCulation (ADCIRC) model [10, 11] which has wide usage in both academia, industry, and government, see, e.g., [12, 13, 14]. ADCIRC solves a surrogate to the SWE in which the continuity equation is replaced by a generalized wave continuity equation [15]. To discretize in space, ADCIRC uses a FE formulation based on the Bubnov-Galerkin, or continuous Galerkin (CG) method [16, 17] and thus continuous polynomial basis functions. As a result of this CG method, element-wise conservation of mass is not guaranteed in ADCIRC simulations. To discretize in time, ADCIRC offers several options, including fully explicit and implicit-explicit flavors of finite differences [11]. ADCIRC has successfully been applied to model compound flooding during Hurricane Harvey (2017) [6, 18] by incorporating rainfall runoff through riverine boundary conditions.
@@
We also mention two other FE based numerical models from recent literature: the Adaptive Hydraulics (AdH) [19] and Semi-implicit Cross-scale Hydro-science Integrated System Model (SCHISM) [20] which both use CG methods in their spatial FE discretizations. While these FE models have the potential to model compound floods as demonstrated in e.g., [21], the conservation properties of the underlying CG method are not exact.
@@
A model that is related to ADCIRC is the discontinuous Galerkin Shallow Water Equation Model (DG-SWEM) [22, 23, 24], which solves the conservative SWE using a local discontinuous Galerkin (DG) method with an explicit strong stability preserving Runge Kutta time stepping scheme. In the recent paper [25], we extended DG-SWEM to incorporate rainfall onto the FE mesh in an effort to model compound floods from rainfall and storm surge. These results indicated that a exact locally mass conserving FE method can accommodate rainfall onto the computational mesh without inducing notable computational issues. The approach in [25] was to modify the right hand side (RHS) of the continuity equation in the SWE with a source. This source term was defined, e.g., from parametric rainfall models from literature such as R-CLIPER.
@@
In this paper, we present a new numerical model based on mixed continuous and discontinuous approximation spaces for the SWE. In particular, we develop a model in which the continuity equation is modified with a nonzero RHS is discretized using a DG scheme and the momentum equation using a CG scheme. Hence, the proposed method inherits the exact local conservation properties of the DG method and the efficiency of the CG method for the momentum equation keeps the solution cost as low as possible. This will allow us to incorporate rainfall onto the FE mesh in a fashion analogous to [25] and thus model compound flooding from runoff and storm surge in a single computational model. Such solution schemes have been investigated in past works as well, e.g., [26]. However, the present work differs in the use of the conservative SWE, as well as its application to large scale computational domains. Last, we also mention the open-source code SWEMniCS [27], where an implicit approach was taken to solve the SWE using a mix of continuous and discontinuous function spaces. The SWEMniCS framework is currently in development but does not presently support large-scale inundation studies with the mixed function space approach.
@@
In the following, we introduce, verify, and validate the proposed numerical model for compound flooding simulations. In section 2, we introduce the governing model, i.e., the modified SWE. Sections 3 - 6 detail the FE discretization of the SWE and present an overview of the implementation details into computer software. Sections 7 - 8 present a comprehensive numerical study of our methodology including small benchmark tests and large scale hurricanes with compound flood effects. Finally, in Section 9, we draw conclusions and discuss potential future research directions.
@@ 2 Governing Equations
The governing two-dimensional shallow water equations (SWE) which consist of the conservative depth-averaged equations of mass conservation as well as $x$ and $y$ momentum conservation [28] are the model used in this work. This set of equations is stated as follows (see Figure 1 for a schematic of the shallow water depths):
@@
where $\zeta$ is the free surface elevation (positive upwards from the geoid), $h_{b}$ the bathymetry (positive downwards from the geoid), $H$ is the total water column, $\mathbf{u}=\{u_{x},u_{y}\}^{\text{T}}$ is the depth-averaged velocity field, $\kappa$ is the bottom friction factor, $R$ is the mass source term, and the source terms $F_{x},F_{y}$ represent other forces. These can include Coriolis force, tidal potential forces, wind stresses, and wave radiation stresses, and vertically integrated lateral stresses [29]. The bottom friction terms can assume a linear or nonlinear form, or a mix of both. For the linear formulation, $\kappa$ is constant; in the nonlinear formulation $\kappa=C_{d}\sqrt{u_{x}^{2}+v_{x}^{2}}$ where $C_{d}$ is the drag coefficient. Commonly used examples of the nonlinear formulation includes Manning’s $n$ [30], and Chezy [31] friction laws. Coriolis force is expressed as $fHu_{y}$ and $fHu_{x}$, where $f=2\times 7.29212\times 10^{-5}\times\sin{\phi}$ and $\phi$ is the latitude. When wind stress is included, it contributes through the pressure gradient terms $\frac{\partial P}{\partial x}$ and $\frac{\partial P}{\partial y}$, acting on the fluid surface. The computational domain is denoted by $\Omega$ and its boundary $\Gamma$ is typically specified by three distinctive sections $\Gamma=\Gamma_{ocean}\cup\Gamma_{land}\cup\Gamma_{river}$. On these sections, the boundary conditions include specified elevation conditions, zero normal flow, and specified normal flow, respectively.
@@ 3.1 Continuity Equation
Our solver uses the Dubiner basis functions [32] for $\psi_{n}$ which are orthogonal and therefore yield a diagonal mass matrix which is trivially inverted. Note that we use the notation $\psi_{n},\;n=1,2,3$ to represent the linear Dubiner bases $\phi_{00},\phi_{01},\phi_{10}$. We will also frequently refer directly to the vector of modal coefficients:
@@
In using the DG method to solve the full SWE, it is standard to use a higher-order total-variation bounded Runge-Kutta scheme like the Strong Stability-Preserving RK scheme (SSPRK) [33]. In this coupled case, we have not encountered stability issues with only using forward Euler, and this significantly reduces the computational cost.
@@ 3.3 Momentum Equation
To solve the momentum equations, we use the existing CG discretization in ADCIRC as described in the theory report [29]. Specifically, we use the (default) nonconservative formulation which transforms Eqs. (1b) and (1c) into
$\frac{\partial u_{x}}{\partial t}$ $=-{u_{x}\frac{\partial u_{x}}{\partial x}}-{u_{y}\frac{\partial u_{x}}{\partial y}}-{g\frac{\partial(\zeta+P_{s}/g\rho_{0})}{\partial x}}+fu_{y}-\frac{\kappa u_{x}}{H\rho_{0}}$ (20)
$\frac{\partial u_{y}}{\partial t}$ $=-{u_{x}\frac{\partial u_{y}}{\partial x}}-{u_{y}\frac{\partial u_{y}}{\partial y}}-{g\frac{\partial(\zeta+P_{s}/g\rho_{0})}{\partial y}}+fu_{x}-\frac{\kappa u_{y}}{H\rho_{0}}$ (21)
where we have limited the forcing terms to be bottom friction and wind forcing for simplicity. Typically we use a quadratic bottom friction $\kappa=C_{d}\sqrt{u_{x}^{2}+u_{y}^{2}}$ where $C_{d}$ is the drag coefficient. We apply the nodal CG spatial discretization, such that each global finite element node $j$ defines a global nodal basis function $\phi_{j}$ which we multiply with (20) and integrate over the whole domain $\Omega$. The goal is then to find $(u_{x},v_{x})\in\mathbf{W}_{h}$ such that:
@@
$\int_{\Omega}{\frac{\partial u_{y}}{\partial t}}{\phi_{j}}$ $=-\int_{\Omega}{u_{x}\frac{\partial V}{\partial x}}{\phi_{j}}-\int_{\Omega}{u_{y}\frac{\partial u_{y}}{\partial y}}{\phi_{j}}-\int_{\Omega}{g\frac{\partial\zeta+P/g\rho}{\partial y}}{\phi_{j}}$
@@
for all $\phi_{i}\in W_{h}$. Applying quadrature rules for linear triangular elements on each term then yields the semidiscrete equations for $\partial u_{x}/\partial t$ and $\partial u_{y}/\partial t$. The time derivative term is approximated using a lumped approach to produce a diagonal mass matrix on the LHS:
$\int_{\Omega}\frac{\partial u_{x}}{\partial t}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{d}{dt}u_{x}(\mathbf{x}_{j}),$ (24)
$\int_{\Omega}\frac{\partial u_{y}}{\partial t}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{d}{dt}u_{y}(\mathbf{x}_{j}).$ (25)
@@
$\int_{\Omega}u_{x}\frac{\partial u_{y}}{\partial x}\phi_{j}\approx\sum_{n\in E_{j}}\frac{A_{n}}{3}\langle{u}_{x}\rangle_{n}\left(\frac{\partial u_{y}}{\partial x}\right)_{n},$ (28)
$\int_{\Omega}u_{y}\frac{\partial u_{y}}{\partial y}\phi_{j}\approx\sum_{n\in E_{j}}\frac{A_{n}}{3}\langle{u}_{y}\rangle_{n}\left(\frac{\partial u_{y}}{\partial y}\right)_{n},$ (29)
@@
$\int_{\Omega}g\frac{\partial\zeta}{\partial x}\phi_{j}$ $\approx\frac{1}{3}g\sum_{n\in E_{j}}A_{n}\left(\frac{\partial\zeta+P/g\rho}{\partial x}\right)_{n},$ (30)
$\int_{\Omega}g\frac{\partial\zeta}{\partial y}\phi_{j}$ $\approx\frac{1}{3}g\sum_{n\in E_{j}}A_{n}\left(\frac{\partial\zeta+P/g\rho}{\partial y}\right)_{n}.$ (31)
@@
$\int_{\Omega}\frac{\kappa u_{x}}{H\rho_{0}}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{\kappa u_{x}}{H\rho_{0}},$ (34)
$\int_{\Omega}\frac{\kappa u_{y}}{H\rho_{0}}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{\kappa u_{y}}{H\rho_{0}}.$ (35)
The resulting semidiscrete equations for each node $j$ are:
@@
$\frac{du_{y}}{dt}(\mathbf{x}_{j})$ $=-\frac{1}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\left[\langle u_{x}\rangle_{n}\left(\frac{\partial u_{y}}{\partial x}\right)_{n}+\langle u_{y}\rangle_{n}\left(\frac{\partial u_{y}}{\partial y}\right)_{n}\right]$
$-\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\left(\frac{\partial[\zeta+P/g\rho]}{\partial y}\right)_{n}-\frac{\Lambda_{j}}{3}fu_{x}(\mathbf{x}_{j})-\frac{\kappa u_{y}(\mathbf{x}_{j})}{H\rho}.$ (37)
ADCIRC utilizes a two-level time discretization, whereby some terms are averaged between current and future time steps $s$ and $s+1$. The LHS terms use a forward difference:
$\frac{du_{x}}{dt}(\mathbf{x}_{j})$ $\approx\frac{u^{s+1}_{x}(\mathbf{x}_{j})-u^{s}_{x}(\mathbf{x}_{j})}{\Delta t}$ (38)
$\frac{du_{y}}{dt}(\mathbf{x}_{j})$ $\approx\frac{u^{s+1}_{y}(\mathbf{x}_{j})-u^{s}_{y}(\mathbf{x}_{j})}{\Delta t},$ (39)
@@
$\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial x}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial x}\right)_{n}$ (40)
$\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial y}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial y}\right)_{n},$ (41)
@@
$\frac{1}{2}f(u_{y}^{s}+u_{y}^{s+1}),$ (43)
and:
$\frac{1}{2}\left(\frac{\kappa u_{x}^{s}}{H^{s}\rho}+\frac{\kappa u_{x}^{s+1}}{H^{s+1}\rho_{0}}\right),$ (44)
$\frac{1}{2}\left(\frac{\kappa u_{y}^{s}}{H^{s}\rho}+\frac{\kappa u_{y}^{s+1}}{H^{s+1}\rho_{0}}\right).$ (45)
We then group the future terms to the LHS and current terms to the RHS. The final discretization is shown below.
@@
$-\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial x}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial x}\right)_{n},$ (46)
$\left(1+\frac{\Delta t\;\kappa}{2H(\mathbf{x}_{j})}\right)$ $u_{y}^{s+1}(\mathbf{x}_{j})-\frac{f\Delta t}{2}u_{x}^{s+1}(\mathbf{x}_{j})=$
$\left(1+\frac{\Delta t\;\kappa}{2H(\mathbf{x}_{j})}\right)u_{y}^{s}(\mathbf{x}_{j})$
$-\frac{1}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\left[\langle u_{x}^{s}\rangle_{n}\left(\frac{\partial u_{y}^{s}}{\partial x}\right)_{n}+\langle u_{y}^{s}\rangle_{n}\left(\frac{\partial u_{y}^{s}}{\partial y}\right)_{n}\right]$
$-\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial y}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial y}\right)_{n}.$ (47)
This yields a $2\times 2$ linear system for $u_{x}^{s+1},u_{y}^{s+1}$ at each node $j$. After abstracting away the known terms, this can be viewed more concisely as

### v2

@@ 1 Introduction
Numerical models based on finite element (FE) discretizations to simulate hydrodynamics in coasts, estuaries, and the ocean have widespread use in engineering, marine, and other scientific disciplines. The model that is the focus of our current investigation are the depth-averaged two dimensional Shallow Water Equations (SWE) [vreugdenhil1994numerical]. In this particular case, we use the SWE to model flows as a result of the interaction between rivers, rainfall runoff, and storm surge, i.e., compound floods [wahl2015increasing]. Famous examples from the last decade includes Hurricane Harvey (2017) [watson2018characterization] Hurricane Irma (2017) [cangialosi2018tropical], and Hurricane Florence (2018) [callaghan2020extreme]. The flow resulting from the interaction of these processes requires careful treatment to ensure accurate modeling and has therefore been frequently studied in recent literature, see, e.g., [loveland2021developing, santiago2019comprehensive, orton2020flood, kumbier2018investigating]. The aim of the present paper is to introduce a new numerical model in which we use a modified form of the SWE to develop to simulate compound flooding during tropical cyclones [wahl2015increasing].
@@
A widely used existing FE based numerical model for the SWE is the ADvanced CIRCulation (ADCIRC) model [luettich1992adcirc, Pringle2020] which has wide usage in both academia, industry, and government, see, e.g., [bunya2010high, blanton2012urgent, funakoshi2012development]. ADCIRC solves a surrogate to the SWE in which the continuity equation is replaced by a generalized wave continuity equation [Lynch1979-ux]. To discretize in space, ADCIRC uses a FE formulation based on the Bubnov-Galerkin, or continuous Galerkin (CG) method [carey1983finite, ern2004theory] and thus continuous polynomial basis functions. As a result of this CG method, element-wise conservation of mass is not guaranteed in ADCIRC simulations. To discretize in time, ADCIRC offers several options, including fully explicit and implicit-explicit flavors of finite differences [Pringle2020]. ADCIRC has successfully been applied to model compound flooding during Hurricane Harvey (2017) [loveland2021developing, pachev2023one] by incorporating rainfall runoff through riverine boundary conditions.
@@
We also mention two other FE based numerical models from recent literature: the Adaptive Hydraulics (AdH) [savant2020theory] and Semi-implicit Cross-scale Hydro-science Integrated System Model (SCHISM) [zhang2014schism] which both use CG methods in their spatial FE discretizations. While these FE models have the potential to model compound floods as demonstrated in e.g., [zhang2020simulating], the conservation properties of the underlying CG method are not exact.
@@
A model that is related to ADCIRC is the discontinuous Galerkin Shallow Water Equation Model (DG-SWEM) [kubatko2006hp, bunya2009wetting, dawson2011discontinuous], which solves the conservative SWE using a local discontinuous Galerkin (DG) method with an explicit strong stability preserving Runge Kutta time stepping scheme. In the recent paper [wichitrnithed2024discontinuous], we extended DG-SWEM to incorporate rainfall onto the FE mesh in an effort to model compound floods from rainfall and storm surge. These results indicated that a exact locally mass conserving FE method can accommodate rainfall onto the computational mesh without inducing notable computational issues. The approach in [wichitrnithed2024discontinuous] was to modify the right hand side (RHS) of the continuity equation in the SWE with a source. This source term was defined, e.g., from parametric rainfall models from literature such as R-CLIPER.
@@
In this paper, we present a new numerical model based on mixed continuous and discontinuous approximation spaces for the SWE. In particular, we develop a model in which the continuity equation is modified with a nonzero RHS is discretized using a DG scheme and the momentum equation using a CG scheme. Hence, the proposed method inherits the exact local conservation properties of the DG method and the efficiency of the CG method for the momentum equation keeps the solution cost as low as possible. This will allow us to incorporate rainfall onto the FE mesh in a fashion analogous to [wichitrnithed2024discontinuous] and thus model compound flooding from runoff and storm surge in a single computational model. Such solution schemes have been investigated in past works as well, e.g., [dawson2002discontinuous]. However, the present work differs in the use of the conservative SWE, as well as its application to large scale computational domains. Last, we also mention the open-source code SWEMniCS [dawson2024swemnics], where an implicit approach was taken to solve the SWE using a mix of continuous and discontinuous function spaces. The SWEMniCS framework is currently in development but does not presently support large-scale inundation studies with the mixed function space approach.
@@
In the following, we introduce, verify, and validate the proposed numerical model for compound flooding simulations. In section 2, we introduce the governing model, i.e., the modified SWE. Section 3 details the FE discretization of the SWE and present an overview of the implementation details. Sections 6 - 7 present a comprehensive numerical study of our methodology including small benchmark tests and large scale hurricanes with compound flood effects. Finally, in Section 8, we draw conclusions and discuss potential future research directions.
@@ 2 Governing Equations
The governing two-dimensional shallow water equations (SWE) which consist of the conservative depth-averaged equations of mass conservation as well as $x$ and $y$ momentum conservation [tan1992shallow] are the model used in this work. This set of equations is stated as follows (see Figure 1 for a schematic of the shallow water depths):
@@
where $\zeta$ is the free surface elevation (positive upwards from the geoid), $h_{b}$ the bathymetry (positive downwards from the geoid), $H$ is the total water column, $\mathbf{u}=\{u_{x},u_{y}\}^{\text{T}}$ is the depth-averaged velocity field, $\kappa$ is the bottom friction factor, $R$ is the mass source term, and the source terms $F_{x},F_{y}$ represent other forces. These can include Coriolis force, tidal potential forces, wind stresses, and wave radiation stresses, and vertically integrated lateral stresses [luettich2004formulation]. The bottom friction terms can assume a linear or nonlinear form, or a mix of both. For the linear formulation, $\kappa$ is constant; in the nonlinear formulation $\kappa=C_{d}\sqrt{u_{x}^{2}+v_{x}^{2}}$ where $C_{d}$ is the drag coefficient. Commonly used examples of the nonlinear formulation includes Manning’s $n$ [Manning1891], and Chezy [Prony1790] friction laws. Coriolis force is expressed as $fHu_{y}$ and $fHu_{x}$, where $f=2\times 7.29212\times 10^{-5}\times\sin{\phi}$ and $\phi$ is the latitude. When wind stress is included, it contributes through the pressure gradient terms $\frac{\partial P}{\partial x}$ and $\frac{\partial P}{\partial y}$, acting on the fluid surface. The computational domain is denoted by $\Omega$ and its boundary $\Gamma$ is typically specified by three distinctive sections $\Gamma=\Gamma_{ocean}\cup\Gamma_{land}\cup\Gamma_{river}$. On these sections, the boundary conditions include specified elevation conditions, zero normal flow, and specified normal flow, respectively.
@@ 3.1 Continuity Equation
Our solver uses the Dubiner basis functions [Dubiner1991-ts] for $\psi_{n}$ which are orthogonal and therefore yield a diagonal mass matrix which is trivially inverted. Note that we use the notation $\psi_{n},\;n=1,2,3$ to represent the linear Dubiner bases $\phi_{00},\phi_{01},\phi_{10}$. We will also frequently refer directly to the vector of modal coefficients:
@@
In using the DG method to solve the full SWE, it is standard to use a higher-order total-variation bounded Runge-Kutta scheme like the Strong Stability-Preserving RK scheme (SSPRK) [Shu1987-ws]. In this coupled case, we have not encountered stability issues with only using forward Euler, and this significantly reduces the computational cost.
@@ 3.3 Momentum Equation
To solve the momentum equations, we use the existing CG discretization in ADCIRC as described in the theory report [luettich2004formulation]. Specifically, we use the (default) nonconservative formulation which transforms Eqs. (1b) and (1c) into
$\frac{\partial u_{x}}{\partial t}$ $=-{u_{x}\frac{\partial u_{x}}{\partial x}}-{u_{y}\frac{\partial u_{x}}{\partial y}}-{g\frac{\partial(\zeta+P/g\rho)}{\partial x}}+fu_{y}-\frac{\kappa u_{x}}{H\rho}$ (20)
$\frac{\partial u_{y}}{\partial t}$ $=-{u_{x}\frac{\partial u_{y}}{\partial x}}-{u_{y}\frac{\partial u_{y}}{\partial y}}-{g\frac{\partial(\zeta+P/g\rho)}{\partial y}}+fu_{x}-\frac{\kappa u_{y}}{H\rho}$ (21)
where we have limited the forcing terms to be bottom friction and wind forcing for simplicity. Here, $f$ represents the Coriolis term, $\rho$ the water density, and $P$ the atmospheric pressure at sea surface. We typically use a quadratic bottom friction $\kappa=C_{d}\sqrt{u_{x}^{2}+u_{y}^{2}}$ where $C_{d}$ is the drag coefficient. We apply the nodal CG spatial discretization, such that each global finite element node $j$ defines a global nodal basis function $\phi_{j}$ which we multiply with (20) and integrate over the whole domain $\Omega$. The goal is then to find $(u_{x},v_{x})\in\mathbf{W}_{h}$ such that:
@@
$\int_{\Omega}{\frac{\partial u_{y}}{\partial t}}{\phi_{j}}$ $=-\int_{\Omega}{u_{x}\frac{\partial u_{y}}{\partial x}}{\phi_{j}}-\int_{\Omega}{u_{y}\frac{\partial u_{y}}{\partial y}}{\phi_{j}}-\int_{\Omega}{g\frac{\partial\zeta+P/g\rho}{\partial y}}{\phi_{j}}$
@@
for all $\phi_{i}\in W_{h}$. Applying quadrature rules for linear triangular elements on each term then yields the semidiscrete equations for $\partial u_{x}/\partial t$ and $\partial u_{y}/\partial t$. The time derivative term is approximated using a lumped approach to produce a diagonal mass matrix on the LHS (in the following we omit the formulation for the $y$-momentum equation for brevity):
$\int_{\Omega}\frac{\partial u_{x}}{\partial t}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{d}{dt}u_{x}(\mathbf{x}_{j}).$ (24)
@@
$\int_{\Omega}g\frac{\partial\zeta}{\partial x}\phi_{j}$ $\approx\frac{1}{3}g\sum_{n\in E_{j}}A_{n}\left(\frac{\partial\zeta+P/g\rho}{\partial x}\right)_{n}.$ (29)
@@
$\int_{\Omega}\frac{\kappa u_{x}}{H\rho_{0}}\phi_{j}$ $\approx\frac{\Lambda_{j}}{3}\frac{\kappa u_{x}}{H\rho_{0}}.$ (32)
The resulting semidiscrete equations for $u_{x}$ at each node $j$ are:
@@
and similarly for $u_{y}$. ADCIRC utilizes a two-level time discretization, whereby some terms are averaged between current and future time steps $s$ and $s+1$. The LHS terms use a forward difference:
$\frac{du_{x}}{dt}(\mathbf{x}_{j})$ $\approx\frac{u^{s+1}_{x}(\mathbf{x}_{j})-u^{s}_{x}(\mathbf{x}_{j})}{\Delta t},$ (36)
@@
$\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial x}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial x}\right)_{n},$ (38)
@@
$\frac{1}{2}\left(\frac{\kappa u_{x}^{s}}{H^{s}\rho}+\frac{\kappa u_{x}^{s+1}}{H^{s+1}\rho_{0}}\right).$ (40)
We then group the future terms to the LHS and current terms to the RHS. The final discretization for $u_{x}$ is shown below.
@@
$-\frac{g}{\Lambda_{j}}\sum_{n\in E_{j}}A_{n}\frac{1}{2}\left(\frac{\partial[\zeta+P/g\rho]^{s}}{\partial x}+\frac{\partial[\zeta+P/g\rho]^{s+1}}{\partial x}\right)_{n}.$ (42)
When coupled with the discretization of $u_{y}$, this yields a $2\times 2$ linear system for $u_{x}^{s+1},u_{y}^{s+1}$ at each node $j$. After abstracting away the known terms, this can be viewed more concisely as